# Inferring Outcome Means of Exponential Family Distributions Estimated by Deep Neural Networks

### v2

@@ Appendix B Proofs in Appendix A
$\mathbb{E}(Y|\eta)=\psi^{\prime}(\eta),\quad\operatorname{{\rm Var}}(Y|\eta)=\psi^{\prime\prime}(\eta)\geq 0.$
We conclude that $\psi(\cdot)$ is a convex function. By Assumption 3.1, the non-degeneracy and the existence of distinct values in the support of $p(y|\eta)=h(y)\exp(\eta y-\psi(\eta))$ ensure that $\operatorname{{\rm Var}}(Y|\eta)>0$ for all $\eta\in[-C,C]$. This guarantees that the second derivative of the log-partition function, $\psi^{\prime\prime}(\eta)=\operatorname{{\rm Var}}(Y|\eta)$, is strictly positive. Furthermore, since $\mathbb{E}(Y|\eta)$ and $\mathbb{E}(Y^{2}|\eta)$ are defined as smooth integrals with respect to $p(y|\eta)$, they are continuous in $\eta$. As a result, $\psi^{\prime\prime}(\eta)$ is also continuous. By the compactness of $\eta\in[-C,C]$, there exists a constant $c>0$ such that:
@@
We propose several lemmas to address the general nonparametric regression cases, serving as the basis for the proof of Proposition A.1. Recall the definition of $R_{n}(\widehat{f}_{n},f_{0})$ in (2.6):
@@
In the following lemmas, we demonstrate that the differences between $R_{n}(\widehat{f}_{n},f_{0})$ and $\widehat{R}_{n}(\widehat{f}_{n},f_{0})$ are small. Some of the proof techniques are inspired by Schmidt-Hieber (2020), but our analysis extends to a more general regression framework, and our proofs differ accordingly. In fact, the settings considered by Schmidt-Hieber (2020) are special cases within our broader context. Additionally, we provide upper and lower bounds for $\widehat{R}_{n}(\widehat{f}_{n},f_{0})$. These results will be used for proving Proposition A.1.
Lemma B.1.
@@
Lemma B.1 establishes the differences between $R_{n}(\widehat{f}_{n},f_{0})$ and $\widehat{R}_{n}(\widehat{f}_{n},f_{0})$. We now propose the following lemma, which addresses the challenges arising from the dependency of $\varepsilon_{i}$ on the covariates $\mathbf{x}_{i}$. This approach is different from previous work, such as Schmidt-Hieber (2020) and Fan and Gu (2024), which assumed independence between $\varepsilon_{i}$ and $\mathbf{x}_{i}$.
Lemma B.2.
Suppose that Assumptions 3.1 and 3.2 hold. Define $\varepsilon_{i}=y_{i}-\psi^{\prime}(f_{0}(\mathbf{x}_{i}))$. For any estimator $\widetilde{f}\in\mathcal{F}$, there exists constant $C>0$ such that
@@
The main difficulty in Lemma B.2 is the dependence of $\varepsilon_{i}$ and $\mathbf{x}_{i}$ under the GNRM framework. Unlike traditional concentration inequalities, here we apply covering number analysis and truncation techniques to address the heteroskedasticity in GNRMs. Details can be refered to Section B.2. With Lemma B.2, we establish the following two lemmas, which provides the upper and lower bound of $\widehat{R}_{n}(\widehat{f}_{n},f_{0})$.
Lemma B.3.
Suppose that Assumptions 3.1 and 3.2 hold. For any fixed $f\in\mathcal{F}$, it holds that
$\widehat{R}_{n}(\widehat{f}_{n},f_{0})\leq\inf_{f\in\mathcal{F}}\mathbb{E}\ell(\bm{X};f,f_{0})+\sqrt{\frac{2C\widehat{R}_{n}(\widetilde{f},f_{0})\log\mathcal{N}_{n}}{n}}+\sqrt{2C}\bigg{(}\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\widetilde{\delta}\bigg{)}$
$+2\widetilde{\delta}C+\Delta_{n}(\widehat{f}_{n},f_{0}).$
@@
Lemma B.4.
Suppose that Assumptions 3.1 and 3.2 hold. It holds that
$\widehat{R}_{n}(\widehat{f}_{n},f_{0})\geq\frac{1}{1+\rho}\cdot\bigg{(}\Delta_{n}(\widehat{f}_{n},f_{0})-2\sqrt{2C}\bigg{(}\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\widetilde{\delta}\bigg{)}-4\widetilde{\delta}C-\frac{C\log\mathcal{N}_{n}}{2n}\bigg{)}$
@@
It is clear to see that $\ell(\mathbf{x};\widehat{f}_{n},f_{0})\geq 0$ by Lemma A.4, and
@@
$|R_{n}(\widehat{f}_{n},f_{0})-\widehat{R}_{n}(\widehat{f}_{n},f_{0})|\leq\mathbb{E}\bigg{\{}\bigg{|}\frac{1}{n}\sum_{i=1}^{n}g_{j^{*}}(\bm{X}^{\prime}_{i},\mathbf{x}_{i})\bigg{|}\bigg{\}}+4\widetilde{\delta}C_{\operatorname{Lip}},$ (B.1)
@@
$\begin{split}g_{j^{*}}(\bm{X}^{\prime}_{i},\mathbf{x}_{i})&=\ell(\bm{X}^{\prime}_{i};f_{j^{*}},f_{0})-\ell(\mathbf{x}_{i};f_{j^{*}},f_{0}).\end{split}$ (B.2)
@@
and it follows from (B.1) and the inequality above that
$|R_{n}(\widehat{f}_{n},f_{0})-\widehat{R}_{n}(\widehat{f}_{n},f_{0})|\leq\mathbb{E}\frac{T}{n}\cdot\Big{(}\sqrt{\log\mathcal{N}_{n}/n}+U+2\widetilde{\delta}C_{\operatorname{Lip}}\Big{)}+4\widetilde{\delta}C_{\operatorname{Lip}}.$ (B.3)
Hence, it remains for us to consider the bound of (B.3). By the definition of $g_{j}(\bm{X}^{\prime}_{i},\mathbf{x}_{i})$, it is clear that $\mathbb{E}g_{j}(\bm{X}^{\prime}_{i},\mathbf{x}_{i})=0$,
@@
$\operatorname{{\rm Var}}\big{(}g_{j}(\bm{X}^{\prime}_{i},\mathbf{x}_{i})\big{)}$ $=2\operatorname{{\rm Var}}(\ell(\bm{X};f_{j},f_{0}))$
@@
$\mathbb{P}(|\sum_{i=1}^{n}\xi_{i}|\geq t_{0})\leq 2\exp\bigg{(}-\frac{t_{0}^{2}}{2Mt_{0}/3+2\sum_{i=1}^{n}\operatorname{{\rm Var}}(\xi_{i})}\bigg{)}.$
@@
Here, we take $\xi_{i}=g_{j}(\bm{X}^{\prime}_{j},\mathbf{x}_{j})/R$, and let $t_{0}=t\max\{\sqrt{\log\mathcal{N}_{n}/n},\mathbb{E}^{1/2}[\ell(\bm{X};f_{j},f_{0})]\}$. According to (B.4), we can bound the value of $\mathbb{E}T$ and $\mathbb{E}T^{2}$, which will further be applied in (B.3). With the assumption above, we have $R\geq 1$, and the equation above can be simplified as
$\mathbb{P}(T/R\geq t)\leq 2\mathcal{N}_{n}\cdot\exp\bigg{(}-\frac{t^{2}}{\frac{4}{3}t\max^{-1}\{\sqrt{\log\mathcal{N}_{n}/n},\mathbb{E}^{1/2}[\ell(\bm{X};f_{j},f_{0})]\}+4n}\bigg{)}.$ (B.4)
@@
[…]
$=6\sqrt{n\log\mathcal{N}_{n}}+4\sqrt{\frac{n}{\log\mathcal{N}_{n}}}.$ (B.5)
Here, the inequality comes from (B.4) that when $t\geq 6\sqrt{n\log\mathcal{N}_{n}}$, $\mathbb{P}(T/R>t)\leq 2\mathcal{N}_{n}\exp\Big{(}-\frac{t\sqrt{\log\mathcal{N}_{n}}}{2\sqrt{n}}\Big{)}$, and the last inequality is by some algebra.
@@
[…]
$=36n\log\mathcal{N}_{n}+16\cdot\frac{1+3\log\mathcal{N}_{n}}{\log\mathcal{N}_{n}}.$ (B.6)
[…]
Combining (B.5), (B.6) with (B.3), and noting that $\mathbb{E}U^{2}=R_{n}(\widehat{f}_{n},f_{0})$ where $U$ is defined below (B.2), we have
@@
Here, the first inequality is by the Cauchy-Schwarz inequality, the second inequality is by (B.5) and (B.6) and the last inequality holds when $n$ is large enough.
B.2 Proof of Lemma B.2
@@
$\leq 2\widetilde{\delta}C+\frac{1}{n}\bigg{|}\mathbb{E}\sum_{i=1}^{n}\varepsilon_{i}f_{j^{\prime}}(\mathbf{x}_{i})\bigg{|}.$ (B.7)
@@
$=\frac{1}{n}\bigg{(}\sqrt{\frac{\log\mathcal{N}_{n}}{n}}+\widehat{R}_{n}^{1/2}(\widetilde{f},f_{0})/\sqrt{c}+\widetilde{\delta}\bigg{)}\cdot\mathbb{E}^{1/2}\frac{\Big{|}\sum_{i=1}^{n}\varepsilon_{i}(f_{j^{\prime}}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i}))\Big{|}^{2}}{r_{j^{\prime}}^{2}}.$ (B.8)
Here, the first equality is by $\mathbb{E}\varepsilon_{i}f_{0}(\mathbf{x}_{i})=\mathbb{E}(\mathbb{E}\varepsilon_{i}f_{0}(\mathbf{x}_{i})|\mathbf{x}_{i})=0$, the first inequality is by $r_{j}\leq\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\sum_{i=1}^{n}(f_{j}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i}))^{2}/n}$, the second inequality is by triangle inequality $\|\mathbf{a}+\mathbf{b}\|\leq\|\mathbf{a}\|+\|\mathbf{b}\|$ and $\|f_{j^{\prime}}-\widetilde{f}\|_{\infty}\leq\widetilde{\delta}$, and the last inequality is by Lemma A.4 that $\ell(\mathbf{x};\widetilde{f},f_{0})\geq c(\widetilde{f}(\mathbf{x})-f_{0}(\mathbf{x}))^{2}$ for some constant $c>0$.
@@
$\xi_{j}=\frac{\big{|}\sum_{i=1}^{n}\varepsilon_{i}(f_{j}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i}))\big{|}}{r_{j}}.$ (B.9)
@@
$\operatorname{{\rm Var}}(\xi_{j}|(\mathbf{x}_{i})_{i})$ $=\frac{\sum_{i=1}^{n}\operatorname{{\rm Var}}\big{(}\varepsilon_{i}(f_{j}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i}))|(\mathbf{x}_{i})_{i}\big{)}}{r_{j}^{2}}$
@@
where the first inequality is by $\operatorname{{\rm Var}}(U|X)\leq\mathbb{E}(U^{2}|X)$ for all random variables $X$ and $U$, and the second inequality is by the definition of $r_{j}=\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\vee\sqrt{\sum_{i=1}^{n}(f_{j}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i}))^{2}/n}$, and $\mathbb{E}\big{(}\varepsilon_{i}^{2}|(\mathbf{x}_{i})_{i}\big{)}\leq 2\kappa^{2}$. Also, letting $a_{ij}=[f_{j}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i})]/r_{j}$, we have
@@
[…]
@@
[…]
@@
By the definition of $\xi_{j}$ in (B.9) and with $\xi_{\max}=\max_{j\in[\mathcal{N}]}\xi_{j}$, we have for any $j^{\prime}$,
@@
Combining the inequality above with (B.8), we have
@@
$\leq\sqrt{\frac{2C\widehat{R}_{n}(\widetilde{f},f_{0})\log\mathcal{N}_{n}}{n}}+\sqrt{2C}\bigg{(}\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\widetilde{\delta}\bigg{)}$ (B.10)
for some constant $C>0$. Combining (B.10) with (B.7) we conclude that for any $\widetilde{f}\in\mathcal{F}$,
@@
for some large constant $C>0$, which completes the proof of Lemma B.2.
B.3 Proof of Lemma B.3
Recall the definition of $\Delta_{n}(\widehat{f}_{n},f_{0})$
$\Delta_{n}(\widehat{f}_{n},f_{0})=\mathbb{E}_{f_{0}}\frac{1}{n}\sum_{i=1}^{n}-y_{i}\widehat{f}_{n}(\mathbf{x}_{i})+\psi(\widehat{f}_{n}(\mathbf{x}_{i}))-\inf_{f\in\mathcal{F}(L,\mathbf{p},s,\mathcal{F})}\frac{1}{n}\sum_{i=1}^{n}-y_{i}f(\mathbf{x}_{i})+\psi(f(\mathbf{x}_{i})).$
@@
$\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}-y_{i}\widehat{f}_{n}(\mathbf{x}_{i})+\psi(\widehat{f}_{n}(\mathbf{x}_{i}))\leq\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}-y_{i}f(\mathbf{x}_{i})+\psi(f(\mathbf{x}_{i}))+\Delta_{n}(\widehat{f}_{n},f_{0}).$ (B.11)
@@
$\widehat{R}_{n}(\widehat{f}_{n},f_{0})$ $\leq\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}\ell(\mathbf{x}_{i};f,f_{0})+\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\widehat{f}_{n}(\mathbf{x}_{i})+\Delta_{n}(\widehat{f}_{n},f_{0})$
$=\mathbb{E}\ell(\bm{X};f,f_{0})+\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\widehat{f}_{n}(\mathbf{x}_{i})+\Delta_{n}(\widehat{f}_{n},f_{0})$
@@
$+\sqrt{2C}\bigg{(}\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\widetilde{\delta}\bigg{)}+2\widetilde{\delta}C+\Delta_{n}(\widehat{f}_{n},f_{0}).$
Here, $\varepsilon_{i}=y_{i}-\psi^{\prime}(f_{0}(\mathbf{x}_{i}))$. The first inequality is from (B.11), the first equality is by $\mathbb{E}\ell(\bm{X};f,f_{0})=\mathbb{E}\ell(\mathbf{x}_{i};f,f_{0})$ for any fixed $f\in\mathcal{F}$, and the last inequality is by Lemma B.2.
B.4 Proof of Lemma B.4

### v3

@@ Appendix D Proofs in Section C
$\mathbb{E}(Y|\eta)=\psi^{\prime}(\eta),\quad\mathrm{Var}(Y|\eta)=\psi^{\prime\prime}(\eta)\geq 0.$
We conclude that $\psi(\cdot)$ is a convex function. By Assumption 3.1, the non-degeneracy and the existence of distinct values in the support of $p(y|\eta)=h(y)\exp(\eta y-\psi(\eta))$ ensure that $\mathrm{Var}(Y|\eta)>0$ for all $\eta\in[-C,C]$. This guarantees that the second derivative of the log-partition function, $\psi^{\prime\prime}(\eta)=\mathrm{Var}(Y|\eta)$, is strictly positive. Furthermore, since $\mathbb{E}(Y|\eta)$ and $\mathbb{E}(Y^{2}|\eta)$ are defined as smooth integrals with respect to $p(y|\eta)$, they are continuous in $\eta$. As a result, $\psi^{\prime\prime}(\eta)$ is also continuous. By the compactness of $\eta\in[-C,C]$, there exists a constant $c>0$ such that:
@@
We propose several lemmas to address the general nonparametric regression cases, serving as the basis for the proof of Proposition C.1. Recall the definition of $R_{n}(\widehat{f}_{n},f_{0})$ in (2.5):
@@
In the following lemmas, we demonstrate that the differences between $R_{n}(\widehat{f}_{n},f_{0})$ and $\widehat{R}_{n}(\widehat{f}_{n},f_{0})$ are small. Some of the proof techniques are inspired by Schmidt-Hieber (2020), but our analysis extends to a more general regression framework, and our proofs differ accordingly. In fact, the settings considered by Schmidt-Hieber (2020) are special cases within our broader context. Additionally, we provide upper and lower bounds for $\widehat{R}_{n}(\widehat{f}_{n},f_{0})$. These results will be used for proving Proposition C.1.
Lemma D.1.
@@
Lemma D.1 establishes the differences between $R_{n}(\widehat{f}_{n},f_{0})$ and $\widehat{R}_{n}(\widehat{f}_{n},f_{0})$. We now propose the following lemma, which addresses the challenges arising from the dependency of $\varepsilon_{i}$ on the covariates $\mathbf{x}_{i}$. This approach is different from previous work, such as Schmidt-Hieber (2020) and Fan & Gu (2024), which assumed independence between $\varepsilon_{i}$ and $\mathbf{x}_{i}$.
Lemma D.2.
Suppose that Assumption 3.1 holds. Define $\varepsilon_{i}=y_{i}-\psi^{\prime}(f_{0}(\mathbf{x}_{i}))$. For any estimator $\widetilde{f}\in\mathcal{F}$, there exists constant $C>0$ such that
@@
The main difficulty in Lemma D.2 is the dependence of $\varepsilon_{i}$ and $\mathbf{x}_{i}$ under the GNRM framework. Unlike traditional concentration inequalities, here we apply covering number analysis and truncation techniques to address the heteroskedasticity in GNRMs. Details can be refered to Section D.2. With Lemma D.2, we establish the following two lemmas, which provides the upper and lower bound of $\widehat{R}_{n}(\widehat{f}_{n},f_{0})$.
Lemma D.3.
Suppose that Assumption 3.1 holds. For any fixed $f\in\mathcal{F}$, it holds that
[…]
$+2\widetilde{\delta}C+\Delta_{n}(\widehat{f}_{n}).$
@@
Lemma D.4.
Suppose that Assumption 3.1 holds. It holds that
$\widehat{R}_{n}(\widehat{f}_{n},f_{0})\geq\frac{1}{1+\rho}\cdot\bigg(\Delta_{n}(\widehat{f}_{n})-2\sqrt{2C}\bigg(\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\widetilde{\delta}\bigg)-4\widetilde{\delta}C-\frac{C\log\mathcal{N}_{n}}{2n}\bigg)$
@@
It is clear to see that $\ell(\mathbf{x};\widehat{f}_{n},f_{0})\geq 0$ by Lemma C.4, and
@@
$|R_{n}(\widehat{f}_{n},f_{0})-\widehat{R}_{n}(\widehat{f}_{n},f_{0})|\leq\mathbb{E}\bigg\{\bigg|\frac{1}{n}\sum_{i=1}^{n}g_{j^{*}}(\bm{X}^{\prime}_{i},\mathbf{x}_{i})\bigg|\bigg\}+4\widetilde{\delta}C_{\operatorname{Lip}},$ (D.1)
@@
$\begin{split}g_{j^{*}}(\bm{X}^{\prime}_{i},\mathbf{x}_{i})&=\ell(\bm{X}^{\prime}_{i};f_{j^{*}},f_{0})-\ell(\mathbf{x}_{i};f_{j^{*}},f_{0}).\end{split}$ (D.2)
@@
and it follows from (D.1) and the inequality above that
$|R_{n}(\widehat{f}_{n},f_{0})-\widehat{R}_{n}(\widehat{f}_{n},f_{0})|\leq\mathbb{E}\frac{T}{n}\cdot\Big(\sqrt{\log\mathcal{N}_{n}/n}+U+2\widetilde{\delta}C_{\operatorname{Lip}}\Big)+4\widetilde{\delta}C_{\operatorname{Lip}}.$ (D.3)
Hence, it remains for us to consider the bound of (D.3). By the definition of $g_{j}(\bm{X}^{\prime}_{i},\mathbf{x}_{i})$, it is clear that $\mathbb{E}g_{j}(\bm{X}^{\prime}_{i},\mathbf{x}_{i})=0$,
@@
$\mathrm{Var}\big(g_{j}(\bm{X}^{\prime}_{i},\mathbf{x}_{i})\big)$ $=2\mathrm{Var}(\ell(\bm{X};f_{j},f_{0}))$
@@
$\mathbb{P}(|\sum_{i=1}^{n}\xi_{i}|\geq t_{0})\leq 2\exp\bigg(-\frac{t_{0}^{2}}{2Mt_{0}/3+2\sum_{i=1}^{n}\mathrm{Var}(\xi_{i})}\bigg).$
@@
Here, we take $\xi_{i}=g_{j}(\bm{X}^{\prime}_{j},\mathbf{x}_{j})/R$, and let $t_{0}=t\max\{\sqrt{\log\mathcal{N}_{n}/n},\mathbb{E}^{1/2}[\ell(\bm{X};f_{j},f_{0})]\}$. According to (D.4), we can bound the value of $\mathbb{E}T$ and $\mathbb{E}T^{2}$, which will further be applied in (D.3). With the assumption above, we have $R\geq 1$, and the equation above can be simplified as
$\mathbb{P}(T/R\geq t)\leq 2\mathcal{N}_{n}\cdot\exp\bigg(-\frac{t^{2}}{\frac{4}{3}t\max^{-1}\{\sqrt{\log\mathcal{N}_{n}/n},\mathbb{E}^{1/2}[\ell(\bm{X};f_{j},f_{0})]\}+4n}\bigg).$ (D.4)
@@
$=6\sqrt{n\log\mathcal{N}_{n}}+\int_{6\sqrt{n\log\mathcal{N}_{n}}}^{+\infty}\mathbb{P}(T\geq t)dt$
$\leq 6\sqrt{n\log\mathcal{N}_{n}}+\int_{6\sqrt{n\log\mathcal{N}_{n}}}^{+\infty}2\mathcal{N}_{n}\exp\bigg(-\frac{t\sqrt{\log\mathcal{N}_{n}}}{2\sqrt{n}}\bigg)dt$
$=6\sqrt{n\log\mathcal{N}_{n}}+4\sqrt{\frac{n}{\log\mathcal{N}_{n}}}.$ (D.5)
Here, the inequality comes from (D.4) that when $t\geq 6\sqrt{n\log\mathcal{N}_{n}}$, $\mathbb{P}(T/R>t)\leq 2\mathcal{N}_{n}\exp\Big(-\frac{t\sqrt{\log\mathcal{N}_{n}}}{2\sqrt{n}}\Big)$, and the last inequality is by some algebra.
@@
$=6^{2}n\log\mathcal{N}_{n}+\int_{6^{2}n\log\mathcal{N}_{n}}^{+\infty}\mathbb{P}(T\geq\sqrt{t})dt$
$\leq 6^{2}n\log\mathcal{N}_{n}+\int_{6^{2}n\log\mathcal{N}_{n}}^{+\infty}2\mathcal{N}_{n}\exp\bigg(-\frac{\sqrt{t}\sqrt{\log\mathcal{N}_{n}}}{2\sqrt{n}}\bigg)dt$
$=36n\log\mathcal{N}_{n}+16\cdot\frac{1+3\log\mathcal{N}_{n}}{\log\mathcal{N}_{n}}.$ (D.6)
Here, the inequality comes from (D.4) that when $t\geq 6\sqrt{n\log\mathcal{N}_{n}}$, $\mathbb{P}(T/R>t)\leq 2\mathcal{N}_{n}\exp\Big(-\frac{t\sqrt{\log\mathcal{N}_{n}}}{(4/3+4/(2R+2))\sqrt{n}}\Big)$, and the last equality is by $\int_{b^{2}}^{+\infty}e^{-\sqrt{t}a}dt=2(ab+1)e^{-ab}/a^{2}$.
Combining (D.5), (D.6) with (D.3), and noting that $\mathbb{E}U^{2}=R_{n}(\widehat{f}_{n},f_{0})$ where $U$ is defined below (D.2), we have
@@
Here, the first inequality is by the Cauchy-Schwarz inequality, the second inequality is by (D.5) and (D.6) and the last inequality holds when $n$ is large enough.
D.2 Proof of Lemma D.2
@@
$\leq 2\widetilde{\delta}C+\frac{1}{n}\bigg|\mathbb{E}\sum_{i=1}^{n}\varepsilon_{i}f_{j^{\prime}}(\mathbf{x}_{i})\bigg|.$ (D.7)
@@
$=\frac{1}{n}\bigg(\sqrt{\frac{\log\mathcal{N}_{n}}{n}}+\widehat{R}_{n}^{1/2}(\widetilde{f},f_{0})/\sqrt{c}+\widetilde{\delta}\bigg)\cdot\mathbb{E}^{1/2}\frac{\Big|\sum_{i=1}^{n}\varepsilon_{i}(f_{j^{\prime}}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i}))\Big|^{2}}{r_{j^{\prime}}^{2}}.$ (D.8)
Here, the first equality is by $\mathbb{E}\varepsilon_{i}f_{0}(\mathbf{x}_{i})=\mathbb{E}(\mathbb{E}\varepsilon_{i}f_{0}(\mathbf{x}_{i})|\mathbf{x}_{i})=0$, the first inequality is by $r_{j}\leq\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\sum_{i=1}^{n}(f_{j}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i}))^{2}/n}$, the second inequality is by triangle inequality $\|\mathbf{a}+\mathbf{b}\|\leq\|\mathbf{a}\|+\|\mathbf{b}\|$ and $\|f_{j^{\prime}}-\widetilde{f}\|_{\infty}\leq\widetilde{\delta}$, and the last inequality is by Lemma C.4 that $\ell(\mathbf{x};\widetilde{f},f_{0})\geq c(\widetilde{f}(\mathbf{x})-f_{0}(\mathbf{x}))^{2}$ for some constant $c>0$.
@@
$\xi_{j}=\frac{\big|\sum_{i=1}^{n}\varepsilon_{i}(f_{j}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i}))\big|}{r_{j}}.$ (D.9)
@@
$\mathrm{Var}(\xi_{j}|(\mathbf{x}_{i})_{i})$ $=\frac{\sum_{i=1}^{n}\mathrm{Var}\big(\varepsilon_{i}(f_{j}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i}))|(\mathbf{x}_{i})_{i}\big)}{r_{j}^{2}}$
@@
where the first inequality is by $\mathrm{Var}(U|X)\leq\mathbb{E}(U^{2}|X)$ for all random variables $X$ and $U$, and the second inequality is by the definition of $r_{j}=\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\vee\sqrt{\sum_{i=1}^{n}(f_{j}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i}))^{2}/n}$, and $\mathbb{E}\big(\varepsilon_{i}^{2}|(\mathbf{x}_{i})_{i}\big)\leq 2\kappa^{2}$. Also, letting $a_{ij}=[f_{j}(\mathbf{x}_{i})-f_{0}(\mathbf{x}_{i})]/r_{j}$, we have
@@
$\mathbb{E}(|\xi_{\max}|^{2}|(\mathbf{x}_{i})_{i})$ $=\int_{0}^{+\infty}\mathbb{P}(|\xi_{\max}|^{2}\geq t|(\mathbf{x}_{i})_{i})dt$
$\leq T+\int_{T}^{+\infty}\mathbb{P}(|\xi_{\max}|\geq\sqrt{t}|(\mathbf{x}_{i})_{i})dt$
$\leq T+\mathcal{N}_{n}\cdot\int_{T}^{+\infty}2\exp\bigg\{-c^{\prime}\min\bigg\{\frac{t}{n\kappa^{2}},\frac{\sqrt{t}r_{j}}{2F\kappa}\bigg\}\bigg\}dt$
$\leq T+\mathcal{N}_{n}\cdot\int_{T}^{+\infty}2\exp\bigg\{-c^{\prime}\min\bigg\{\frac{t}{n\kappa^{2}},\frac{\sqrt{t\log\mathcal{N}_{n}}}{2F\kappa\sqrt{n}}\bigg\}\bigg\}dt.$
@@
$\mathbb{E}(|\xi_{\max}|^{2}|(\mathbf{x}_{i})_{i})\leq Cn\log\mathcal{N}_{n}+\mathcal{N}_{n}\cdot\int_{Cn\log\mathcal{N}_{n}}^{+\infty}2\exp\bigg\{-c^{\prime}\frac{\sqrt{t\log\mathcal{N}_{n}}}{2F\kappa\sqrt{n}}\bigg\}dt.$
Again by utilizing the equality $\int_{b^{2}}^{+\infty}e^{-\sqrt{t}a}dt=2(ab+1)e^{-ab}/a^{2}$ we conclude that as long as $C\geq 4F^{2}\kappa^{2}/c^{\prime 2}$,
$\mathbb{E}(|\xi_{\max}|^{2}|(\mathbf{x}_{i})_{i})$ $\leq Cn\log\mathcal{N}_{n}+\mathcal{N}_{n}\cdot\int_{Cn\log\mathcal{N}_{n}}^{+\infty}2\exp\bigg\{-c^{\prime}\frac{\sqrt{t\log\mathcal{N}_{n}}}{2F\kappa\sqrt{n}}\bigg\}dt$
@@
By the definition of $\xi_{j}$ in (D.9) and with $\xi_{\max}=\max_{j\in[\mathcal{N}]}\xi_{j}$, we have for any $j^{\prime}$,
@@
Combining the inequality above with (D.8), we have
@@
[…]
for some constant $C>0$. Combining (D.10) with (D.7) we conclude that for any $\widetilde{f}\in\mathcal{F}$,
@@
for some large constant $C>0$, which completes the proof of Lemma D.2.
D.3 Proof of Lemma D.3
Recall the definition of $\Delta_{n}(\widehat{f}_{n})$
$\Delta_{n}(\widehat{f}_{n})=\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}-y_{i}\widehat{f}_{n}(\mathbf{x}_{i})+\psi(\widehat{f}_{n}(\mathbf{x}_{i}))-\inf_{f\in\mathcal{F}(L,\mathbf{p},s,\mathcal{F})}\frac{1}{n}\sum_{i=1}^{n}-y_{i}f(\mathbf{x}_{i})+\psi(f(\mathbf{x}_{i})).$
@@
$\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}-y_{i}\widehat{f}_{n}(\mathbf{x}_{i})+\psi(\widehat{f}_{n}(\mathbf{x}_{i}))\leq\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}-y_{i}f(\mathbf{x}_{i})+\psi(f(\mathbf{x}_{i}))+\Delta_{n}(\widehat{f}_{n}).$ (D.11)
@@
$\widehat{R}_{n}(\widehat{f}_{n},f_{0})$ $\leq\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}\ell(\mathbf{x}_{i};f,f_{0})+\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\widehat{f}_{n}(\mathbf{x}_{i})+\Delta_{n}(\widehat{f}_{n})$
$=\mathbb{E}\ell(\bm{X};f,f_{0})+\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}\widehat{f}_{n}(\mathbf{x}_{i})+\Delta_{n}(\widehat{f}_{n})$
@@
$+\sqrt{2C}\bigg(\frac{\log\mathcal{N}_{n}}{n}+\sqrt{\frac{\log\mathcal{N}_{n}}{n}}\widetilde{\delta}\bigg)+2\widetilde{\delta}C+\Delta_{n}(\widehat{f}_{n}).$
Here, $\varepsilon_{i}=y_{i}-\psi^{\prime}(f_{0}(\mathbf{x}_{i}))$. The first inequality is from (D.11), the first equality is by $\mathbb{E}\ell(\bm{X};f,f_{0})=\mathbb{E}\ell(\mathbf{x}_{i};f,f_{0})$ for any fixed $f\in\mathcal{F}$, and the last inequality is by Lemma D.2.
D.4 Proof of Lemma D.4